Ammonia contains 82.65% nitrogen and 17.65% Q.hydrogen.if the law of constant proportions is ttue,then the mass of nitrogen required to give 10g ammonia will be

Grade:11

1 Answers

Khimraj
5 years ago
 
The chemical eqn. would be:
N2  +  H2  ------------>  2NH3
                                 10/17 moles
moles of N2 = 5/17
mass of N2 = 5*28/17 = 8.235 gm Nitrogen